# Rule-based Data Selection for Large Language Models

### v2

@@ Abstract
The quality of training data significantly impacts the performance of large language models (LLMs). There are increasing studies using LLMs to rate and select data based on several human-crafted metrics (rules). However, these conventional rule-based approaches often depend too heavily on human heuristics, lack effective metrics for assessing rules, and exhibit limited adaptability to new tasks. In our study, we introduce an innovative rule-based framework that utilizes the orthogonality of score vectors associated with rules as a novel metric for rule evaluations. Our approach includes an automated pipeline that first uses LLMs to generate a diverse set of rules, encompassing various rating dimensions to evaluate data quality. Then it rates a batch of data based on these rules and uses the determinantal point process (DPP) from random matrix theory to select the most orthogonal score vectors, thereby identifying a set of independent rules. These rules are subsequently used to evaluate all data, selecting samples with the highest average scores for downstream tasks such as LLM training. We verify the effectiveness of our method through two experimental setups: 1) comparisons with ground truth ratings and 2) benchmarking LLMs trained with the chosen data. Our comprehensive experiments cover a range of scenarios, including general pre-training and domain-specific fine-tuning in areas such as IMDB, Medical, Math, and Code. The outcomes demonstrate that our DPP-based rule rating method consistently outperforms other approaches, including rule-free rating, uniform sampling, importance resampling, and QuRating, in terms of both rating precision and model performance.
@@ 1 Introduction
Large language models (LLMs) have been widely utilized across a diverse range of applications. Pretraining and fine-tuning these models typically require large and diverse datasets. Studies have found that data quality is critical for training good LLMs (Brown, 2020; Chowdhery et al., 2023; Du et al., 2022; Dubey et al., 2024; Wenzek et al., 2019). For instance, Meta’s LIMA paper (Zhou et al., 2024) demonstrated that using only 1000 carefully curated data samples can achieve better performance than using the original 50k samples. Similar phenomena have been observed in other studies where selecting a subset of high-quality datasets increases the training convergence and model performance (Cao et al., 2023; Hsieh et al., 2023; Xie et al., 2024; Sachdeva et al., 2024; Zhang et al., 2023; Javaheripi et al., 2023).
Recent studies now adopt an approach that employs LLM-as-a-judge to grade data quality according to a set of designed metrics (which we call rules) (Yuan et al., 2024; Wettig et al., 2024; Bai et al., 2022; Mu et al., ). For example, Wettig et al. (2024) rates the pre-training data using LLMs according to four predefined rules. RedPajama (Together AI, 2023) is continuously developing a pool of rules for users to select from, which currently contains over 40 basic criteria that LLM data should satisfy. On specific aspects such as safety, Constitutional AI (Bai et al., 2022) proposed their “constitution”—a set of standard safety criteria—to generate safe synthetic data, and in Huang et al. (2024) they have developed 133 rules. Most recently, OpenAI’s Rule-based Rewarding (Mu et al., ) proposed 21 general safety rules and injected them into the RLHF (reinforcement learning with human feedback) process. This rule-based rating provides greater explainability of data quality and breaks down the challenge of assigning a data point one overall quality score into a simpler task of giving several rule-specific scores. Evidence suggests that this fine-grained approach also yields more accurate rating outcomes (Yuan et al., 2024; Wettig et al., 2024; Bai et al., 2022; Mu et al., ).
@@
Nonetheless, there are several critical problems and challenges. First, designing an effective set of rules is quite difficult, a fact acknowledged by most of the papers above. Current designs in these papers all rely heavily on human heuristics and are sometimes too broad for effective rating. Second, as far as we know, the metrics to evaluate rules are lacking, and there has been no systematic exploration of how different rule choices and sizes impact the outcomes. In previous experiments Bai et al. (2022); Wettig et al. (2024); Together AI (2023), a subset of rules is selected (typically randomly) during the rating process. This selection can significantly influence the rating results and consequently the quality of the sampled data. Furthermore, the utility and impact of rules can vary significantly, and some rules are strongly correlated with each other, as highlighted in Wettig et al. (2024), which introduces redundancy and bias into the rating procedure. Therefore, a critical question arises: with a “constitution” (a pool of rules) in hand, exactly which “laws” (a subset of task-related rules) should be applied to a specific task? Random selection as in Bai et al. (2022) may not be the optimal strategy. A third major drawback is the inflexibility of these rules; they are often designed for specific settings, such as pre-training or safety tasks, not generally applicable across different settings.
@@
In our work, we aim to address these challenges. First, we leverage an LLM (GPT-4 Achiam et al. (2023)) for automatic rule generation, where we include the descriptions of the task and source dataset in the prompt. At this stage, some generated rules are found to be repetitive or redundant, similar to the issues in human-designed rules. Our strategy is to first generate a comprehensive set of rules to ensure a broad diversity that covers various aspects we seek to evaluate in the data, and then filter out repetitive rules. Hence the second step of our approach is to select a subset of rules that are relatively uncorrelated/independent. This is achieved by first using the rules to rate a batch of data, creating one score vector for each rule, and then assessing the independence of rule subsets through the overall orthogonality of their corresponding score vectors. We propose the formula in Section 3 to measure the orthogonality and use determinant point process (DPP) sampling (Macchi, 1975; Borodin & Olshanski, 2000) to identify a subset of independent rules. Once the rules are determined, the third step is to use them to rate all source data and select the high-quality ones. Combining these steps of rule generation, rule-based rating, rule selection by DPP, and data selection, our method establishes a fully automated framework for rule-based data selection (illustrated in Figure 1). Notably, we are the first to introduce the mathematical rule evaluation metric, based on the orthogonality of their score vectors. Moreover, our pipeline does not need human intervention in designing and selecting rules at all. For every new task, one can use the pipeline to get a set of high-quality, task-specific rules at low cost. These address the challenges of existing methods mentioned above. Another advantage of our method is its natural extension, allowing for customization such as re-weighting particular rules, and this is only feasible when the selected rules are relatively “orthonormal”.
Note that our data selection methodology is highly versatile and applicable to a variety of scenarios, including LLM pre-training, fine-tuning on specific domains, RLHF preference data, etc. We conduct experiments to cover a range of tasks and datasets, including general pre-training data and domain-specific data in four domains: IMDB, Medical, Math, and Code. [^1] Code of our experiments: https://anonymous.4open.science/r/DataSelection-F118/ We show that our rule-based data selection typically yields more accurate rating results, thereby enhancing data quality and leading to better performance of the LLM trained with the data. Here is a list of the main contributions of our work:
@@
    Rule-free vs. Rule-based Rating. Our systematic experiments demonstrate that fine-grained rule-based rating outperforms rule-free methods, producing more precise data quality assessments, leading to improved benchmark performance of LLMs.
@@
    Rule Evaluation Metric: We introduce a novel rule evaluation metric designed to promote low correlation and high diversity among rules. We propose the method of using DPP on task-aware rule-rating vectors to select a subset of independent rules.
@@
    Automated Rule-based Rating and Selection Pipeline. We confirm that LLMs are effective rule generators, eliminating the need for manual rule crafting. Our automated pipeline generates the rules, selects the rules, and then chooses data according to rule-based ratings. This entire process operates independently of human heuristics and is free from human biases.
@@
    Cross-Domain and Cross-Model. We validate our method through two approaches: A) comparison with ground truth ratings, and B) training LLMs with selected data and assessing performance across various benchmarks. Our experiments span multiple models, including Pythia-1B and Llama3-8B (fine-tuned with LoRA), and cover diverse domains such as IMDB, Medical, Math, and Code, confirming the versatility and model independence of our approach.
@@ 2 Related Work
LLM data selection. There are different genres of data selection approaches for LLMs. Basic filterings, such as setting thresholds on word lengths, are used in many studies to eliminate low-quality data (Soldaini et al., 2024; Wenzek et al., 2019; Raffel et al., 2020; Conneau & Lample, 2019; Penedo et al., 2023; Laurençon et al., 2022). Fuzzy deduplication is another approach which removes repetitive or similar data samples (Allamanis, 2019; Lee et al., 2021; Abbas et al., 2023; Gao et al., 2020; Jiang et al., 2022). Another method is “heuristic classification”, selecting data based on a predefined quality score, typically measured by similarity to formal sources such as Wikipedia or other human-generated, high-quality datasets (Brown, 2020; Touvron et al., 2023; Chowdhery et al., 2023; Du et al., 2022; Gao et al., 2020; Wenzek et al., 2019). In contrast to this, directly querying LLMs to rate data and use the scores as the quality indicator has become a standard practice in many studies (Li et al., 2023a; Chen et al., 2023; Bai et al., 2022; Wettig et al., 2024; Yuan et al., 2024; Dubois et al., 2024; Li et al., 2023b; Fernandes et al., 2023).
Rule-based rating. There are studies adopting a more fine-grained approach to data quality, distilling it into a finite set of metrics which we refer to as “rules”. For instance, RedPajama (Together AI, 2023) provides over 40 quality rules that serve as basic quality metrics for the users to choose from. More pertinent to our research, there are papers that apply this rule-based idea to rate LLM data. For example, Yuan et al. (2024) assigns a score out of 5 to each data point, awarding 1 point for each of the 5 predefined criteria met. In Wettig et al. (2024), the authors designed four general rules to rate and select data for LLM pre-training. (Sun et al., 2024) proposed 16 human-crafted rules to evaluate the desirable quality of response data. The rule-based approach is also utilized in more targeted applications, such as ensuring data safety. Constitutional AI designed 16 general safety critique rules to revise synthetic data, enhancing data safety (Bai et al., 2022). This revision process involves iterative steps where a random subset of rules from the “constitution” (the entire set of rules) is applied. Additionally, in Mu et al. , the score generated by an LLM grader according to a set of 21 safety rules is integrated directly into the RLHF process as an additional reward. In Wang et al. (2024), they design a composite reward model in RLHF, trained using rule-based ratings. As noted earlier in the introduction, the rules employed in the literature exhibit several critical issues. They often depend heavily on human heuristics for design, lack robust rule evaluation metrics and exploration of rule sizes, and demonstrate limited versatility for new tasks or for customization. Our goal is to address these challenges using our proposed framework.
@@ 3.1 Definitions and Notations
  • $\mathcal{B}\subseteq\mathcal{D}$: a batch of data samples, randomly selected for evaluating the correlation of rules during the rule selection step, with its size denoted by $n\stackrel{{\scriptstyle\text{def}}}{{=}}|\mathcal{B}|$.
@@
Measure orthogonality: We propose a metric for selecting rules based on the orthogonality of score vectors. Here we introduce a mathematical definition to quantify the orthogonality or correlation of a set of score vectors. Given a score matrix $\bar{{\bm{S}}}\in\mathbb{R}^{n\times r}$ such that the columns are the score vectors of dimensions $n$. We begin by computing the covariance matrix $Cov(\bar{{\bm{S}}})$ for the columns of $\bar{{\bm{S}}}$, whose entries are defined by
$Cov(\bar{{\bm{S}}})_{i,j}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{n}\sum_{k=1}^{n}(S_{k,i}-\mu_{i})(S_{k,j}-\mu_{j}),\qquad 1\leq i,j\leq r,$
where each $\mu_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{n}\sum_{k=1}^{n}S_{k,i}$ is the sample mean for rule $i$. Then define the sample correlation matrix as $Corr(\bar{{\bm{S}}})\in\mathbb{R}^{r\times r}$ where
$Corr(\bar{{\bm{S}}})_{i,j}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{Cov(\bar{{\bm{S}}})_{i,j}}{\sqrt{Cov(\bar{{\bm{S}}})_{i,i}\cdot Cov(\bar{{\bm{S}}})_{j,j}}},\qquad 1\leq i,j\leq r.$
Commonly used libraries such as Numpy provide straightforward functions to compute the correlation matrix. We introduce the concept of rule correlation, which quantifies the degree of correlation/dependence for a given rating submatrix $\bar{{\bm{S}}}$, defined as follows:
$\rho(\bar{{\bm{S}}})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{r}\|Corr(\bar{{\bm{S}}})-{\bm{I}}_{r}\|_{F}=\frac{1}{r}\sqrt{\sum_{i\neq j}Corr(\bar{{\bm{S}}})_{i,j}^{2}}.$ (1)
Here ${\bm{I}}_{r}\in\mathbb{R}^{r\times r}$ is the identity matrix, and $\|\cdot\|_{F}$ represents the Frobenius norm. This metric quantifies how much the columns of $\bar{{\bm{S}}}$ deviate from orthogonality, by measuring the deviation of its correlation matrix from the identity matrix. The second equality in 1 provides another intuitive understanding: $\rho(\bar{{\bm{S}}})$ essentially aggregates the correlations of all pairwise correlations of rules $(i,j)$ for $i\neq j$.

### v3

# Selection of LLM Fine-Tuning Data based on Orthogonal Rules

@@ Abstract
High-quality training data is critical to the performance of large language models (LLMs). Recent work has explored using LLMs to rate and select data based on a small set of human-designed criteria (rules), but these approaches often rely heavily on heuristics, lack principled metrics for rule evaluation, and generalize poorly to new tasks. We propose a novel rule-based data selection framework that introduces a metric based on the orthogonality of rule score vectors to evaluate and select complementary rules. Our automated pipeline first uses LLMs to generate diverse rules covering multiple aspects of data quality, then rates samples according to these rules and applies the determinantal point process (DPP) to select the most independent rules. These rules are then used to score the full dataset, and high-scoring samples are selected for downstream tasks such as LLM fine-tuning. We evaluate our framework in two experiment setups: (1) alignment with ground-truth ratings and (2) performance of LLMs fine-tuned on the selected data. Experiments across IMDB, Medical, Math, and Code domains demonstrate that our DPP-based rule selection consistently improves both rating accuracy and downstream model performance over strong baselines.
Code & Datahttps://github.com/XiaominLi1998/Submission-OrthoRules/
@@ 1 Introduction
Large language models (LLMs) have been widely adopted across a diverse range of applications. Training these models—both during pre-training and fine-tuning—typically requires large and varied datasets. Prior work has shown that data quality plays a critical role in the effectiveness of LLM training (brown2020language; chowdhery2023palm; du2022glam; dubey2024llama; wenzek2019ccnet). For example, Meta’s LIMA paper (zhou2024lima) demonstrated that just 1K carefully curated samples can outperform a much larger set of 50K original samples. Similar findings have emerged from other studies, where selecting high-quality data subsets improves both training efficiency and model performance (cao2023instruction; hsieh2023distilling; xie2024data; sachdeva2024train; zhang2023opt).
A recent trend involves using LLM-as-a-judge to rate data quality based on a set of human-designed metrics, referred to here as “rules” (yuan2024self; wettig2024qurating; bai2022constitutional; mu2024rule). For instance, wettig2024qurating used LLMs to score pre-training data according to four predefined rules. RedPajama (RedPajama2023) built a rule set with over 40 criteria for evaluating LLM training data. In specialized domains like safety, Constitutional AI (bai2022constitutional) defined a set of “constitutions” for generating safe synthetic data; huang2024collective later expanded this to 133 rules. OpenAI’s Rule-based Rewarding (mu2024rule) similarly introduced 21 general safety rules as part of the reinforcement learning from human feedback (RLHF) pipeline. These rule-based approaches provide better interpretability by assigning granular, rule-specific scores rather than a single opaque quality label, and studies have shown this fine-grained strategy leads to more accurate assessments (yuan2024self; wettig2024qurating; bai2022constitutional; mu2024rule).
@@
Despite this progress, several challenges remain. First, designing an effective rule set is difficult, as acknowledged in the prior work mentioned above. Most current rules rely heavily on human heuristics and are often too broad to yield useful signal. Second, there is a lack of principled metrics for evaluating rules, and little understanding of how rule quality and quantity affect downstream outcomes. Prior work (bai2022constitutional; wettig2024qurating; RedPajama2023) typically randomly selects a subset of rules for scoring, a decision that can substantially impact the resulting data quality. Additionally, many rules are correlated, introducing redundancy and potential bias in the ratings (wettig2024qurating). This raises an important question: with a “constitution” (a pool of rules) in hand, exactly which “laws” (a subset of task-related rules) should be applied to a specific task? Random selection as in bai2022constitutional may not be the optimal strategy. A third major drawback is the limited flexibility of these rules; they are often designed for specific settings, such as safety tasks, and lack general applicability across diverse settings.
@@
In our work, we aim to address these challenges. We first leverage GPT-4 (achiam2023gpt) to automatically generate candidate rules, prompting it with descriptions of both the target task and the source dataset. The generated rules are then manually reviewed to ensure clarity and validity. At this stage, some generated rules are found to be repetitive or redundant. Our strategy is to first generate a rule set that can comprehensively cover various data quality aspects, which may include many correlated and redundant rules. The second step is to filter out repetitive ones, and select a subset of rules that are relatively uncorrelated/independent. This is achieved by first using the rules to rate a batch of data, creating one score vector for each rule, and then assessing the independence of rule subsets through the overall orthogonality of their corresponding score vectors. We propose the formula in Section 3 to measure the orthogonality and use determinant point process (DPP) sampling (macchi1975coincidence; borodin2000distributions) to identify a subset of independent rules. Once the rules are determined, the third step is to apply them to rate the full dataset and select the high-quality subset. Combining rule generation, rule-based rating, rule selection by DPP, and data selection, our method establishes a fully automated framework for rule-based data selection (illustrated in Figure 1). To our knowledge, we are the first to introduce a mathematical evaluation metric for rules based on score vector orthogonality. Moreover, this pipeline can be applied to new tasks with minimal manual effort, addressing the shortcomings of existing methods.
We validate our approach across four domains: IMDB, Medical, Math, and Code, by fine-tuning LLMs on data selected by our framework. We show that our rule-based method consistently improves data rating accuracy and leads to better model performance. Below is a list of main contributions of our work:
@@
    Rule-free vs. Rule-based Rating. We provide the first systematic comparison showing that fine-grained, rule-based evaluation outperforms rule-free methods in data quality assessment.
@@
    Rule Evaluation Metric: We introduce a novel rule evaluation metric designed to promote low correlation and high diversity among rules, the first rule evaluation metric for rule-based LLM-as-a-judge to our knowledge. We also propose the method of using DPP on task-aware score vectors to select a subset of independent rules.
@@
    Automated Rule-based Selection Pipeline. We confirm that LLMs are effective rule generators, eliminating the need for manual rule crafting. Our automated pipeline generates the rules, selects the rules, and then uses them to identify high-quality data.
@@
    Cross-Domain and Cross-Model. We validate our method both by comparing against ground-truth ratings and by benchmarking LLMs trained on selected data across multiple domains (IMDB, Medical, Math, Code) and model families (Pythia-1B and Llama3-8B with LoRA), confirming the generality of our approach.
@@ 2 Related Work
Rule-based rating. Some studies adopt a more fine-grained approach to data quality by distilling it into a finite set of metrics, which we refer to as “rules”. For instance, RedPajama (RedPajama2023) provides over 40 quality rules as basic filtering criteria. More relevant to our research, several studies apply this rule-based idea to rate LLM data. yuan2024self rates each sample from 1 to 5 based on how many of five predefined criteria it satisfies. wettig2024qurating uses four general rules to rate and select pre-training data, while sun2024principle proposes 16 handcrafted rules to assess response quality. Rule-based evaluation is also used in safety applications. Constitutional AI (bai2022constitutional) applies a random subset of 16 safety critique rules (called “constitutions”) to iteratively revise synthetic data. In mu2024rule, rule-based scores from 21 safety rules are directly integrated into the RLHF reward, while wang2024helpsteer2 trains a composite reward model using rule-based ratings. As noted earlier, existing rule designs often rely heavily on human heuristics, lack principled evaluation metrics and exploration of rule sizes, and are not easily adaptable. Our framework addresses these limitations.
LLM data selection. There are different genres of data selection approaches for LLMs. Basic filtering, such as setting thresholds on word length, is used in many studies to eliminate low-quality data (soldaini2024dolma; wenzek2019ccnet; raffel2020exploring; penedo2023refinedweb). Fuzzy deduplication removes repetitive or similar samples (allamanis2019adverse; lee2021deduplicating; gao2020pile; jiang2022fuzzydedup). Another method is heuristic classification, selecting data based on similarity to high-quality sources such as Wikipedia (brown2020language; touvron2023llama; chowdhery2023palm; du2022glam; gao2020pile; wenzek2019ccnet). More recently, querying LLMs to rate data directly has become a standard practice (li2023quantity; chen2023alpagasus; bai2022constitutional; wettig2024qurating; yuan2024self; dubois2024alpacafarm). Other methods include coreset and optimization-based subset selection (xia2022moderate; xiao2024feature; borsos2020coresets).
@@ 3.1 Definitions and Notations
  • $\mathcal{B}\subseteq\mathcal{D}$: a batch of data samples, randomly selected for evaluating the correlation of rules during the rule selection step, with its size denoted by $n\stackrel{{\scriptstyle\text{def}}}{{=}}|\mathcal{B}|$ (we use $n=10^{4}$ in experiments in Section 5).
@@
Measure orthogonality: To guide rule selection, we propose a metric based on the orthogonality of score vectors. To achieve this, we introduce a mathematical framework to quantify the degree of orthogonality or correlation among a given set of score vectors. Given $\bar{{\bm{S}}}\in\mathbb{R}^{n\times r}$, define its sample mean as $\mu_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{n}\sum_{k=1}^{n}S_{k,i}$, and sample covariance matrix $\widehat{{\bm{\Sigma}}}(\bar{{\bm{S}}})\in\mathbb{R}^{r\times r}$ by
$\widehat{{\bm{\Sigma}}}_{i,j}(\bar{{\bm{S}}})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{n}\sum_{k=1}^{n}\left(S_{k,i}-\mu_{i}\right)\left(S_{k,j}-\mu_{j}\right).$
Then define the sample correlation matrix $\widehat{{\bm{C}}}(\bar{{\bm{S}}})\in\mathbb{R}^{r\times r}$ as
$\widehat{{\bm{C}}}_{i,j}(\bar{{\bm{S}}})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{\widehat{{\bm{\Sigma}}}_{i,j}(\bar{{\bm{S}}})}{\sqrt{\widehat{{\bm{\Sigma}}}_{i,i}(\bar{{\bm{S}}})\;\widehat{{\bm{\Sigma}}}_{j,j}(\bar{{\bm{S}}})}},\quad 1\leq i,j\leq r.$
To quantify the degree of correlation/dependence for a given rating submatrix $\bar{{\bm{S}}}$, we introduce the quantity called rule correlation:
$\rho(\bar{{\bm{S}}})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{r}\left\|\widehat{{\bm{C}}}(\bar{{\bm{S}}})-{\bm{I}}_{r}\right\|_{F}=\frac{1}{r}\sqrt{\sum_{i\neq j}\left(\widehat{{\bm{C}}}_{i,j}(\bar{{\bm{S}}})\right)^{2}}.$ (1)
Here, ${\bm{I}}_{r}$ is the identity matrix and $\|\cdot\|_{F}$ is the Frobenius norm. Intuitively, $\widehat{{\bm{C}}}(\bar{{\bm{S}}})\approx{\bm{I}}_{r}$ means the columns of $\bar{{\bm{S}}}$ have low pairwise correlations, making $\rho(\bar{{\bm{S}}})$ small. This metric quantifies how much the columns of $\bar{{\bm{S}}}$ deviate from orthogonality, by measuring the deviation of its correlation matrix from the identity matrix. The second equality in (1) provides another intuitive understanding: $\rho(\bar{{\bm{S}}})$ essentially aggregates the correlations of all pairwise correlations of rules $(i,j)$ for $i\neq j$.
To validate our approach of using $\mathcal{B}$, batch of random $n$ samples, to generate rating scores and evaluate rule correlations based on these score vectors, we present the following theorem, which characterizes the concentration of the sample rule correlation around the true rule correlation.
Theorem 1.
Let ${\bm{C}}\in\mathbb{R}^{r\times r}$ be the true correlation matrix among the $r$ rules, i.e.,
${\bm{C}}_{j,\ell}=\frac{\Sigma_{j,\ell}}{\sqrt{\Sigma_{j,j}\Sigma_{\ell,\ell}}},$ (2)
and assume each candidate rule has nontrivial variance ($\Sigma_{j,j}\geq\sigma_{\min}^{2}>0$ for some constant $\sigma_{\min}$). We draw $n$ i.i.d. samples $\{{\bm{x}}^{(k)}\}_{1\leq k\leq n}$ where each ${\bm{x}}^{(k)}\in[0,1]^{r}$ represents ratings for the $k$-th sample based on $r$ rules. From these we form the empirical correlation matrix $\widehat{{\bm{C}}}$. Then there exists a universal constant $c>0$ such that for any $\delta>0$ and sufficiently large $n$,
$\mathbb{P}\left(\left|\rho(\bar{{\bm{S}}})-\frac{1}{r}\left\|{\bm{C}}-\mathbf{I}_{r}\right\|_{F}\right|\leq c\sqrt{\frac{\ln\left(\frac{r^{2}}{\delta}\right)}{n}}\right)>1-\delta.$
In particular, if $\bm{C}$} is close to $\mathbf{I}_{r}$ in Frobenius norm (i.e., if the rules are nearly uncorrelated “in truth”), then the observed rule correlation $\rho(\bar{{\bm{S}}})$ also remains close to zero for sufficiently large $n$.
Proof.
See Appendix A. ∎